OpenStudy (anonymous):

all values of n such that n^4 - 51n^2 + 225 is a prime number with n is integer number ?

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

my job : n^4 - 51n^2 + 225 = p n^4 - 30n^2 + 225 - 21n^2 = p (n^2 - 15)^2 - (sqrt(21)n)^2 = p (n^2 - 15 + sqrt(21)n) ((n^2 - 15 - sqrt(21)n) = p one of the faktor must be equal 1 n^2 - 15 + sqrt(21)n = 1 or n^2 - 15 - sqrt(21)n = 1 like that ?

OpenStudy (anonymous):

and likes there is no n which integer satisfied

OpenStudy (rational):

im thinking of factoring the given expression directly

OpenStudy (rational):

Can you check if below is true : \[ n^4 - 51n^2 + 225 = (n^2-9 n+15) (n^2+9 n+15)\]

OpenStudy (anonymous):

ah, yes that's correct.. how you factor out that ?

OpenStudy (rational):

rest should be easy..

OpenStudy (rational):

for \(M = ab\) to be prime, we must have : 1) \(a=1\) and \(b=p\) or 2) \(a=p\) and \(b=1\)

OpenStudy (anonymous):

ok, i get it n^2 + 9n + 15 = 1 n^2 + 9n + 14 = 0 (n + 2)(n + 7) = 0 n = -2 or n = -7 and the rest there is no n which integer

OpenStudy (anonymous):

thank you very much for your help :D

OpenStudy (rational):

hey so what are the "n" values that make the given expression prime ?